### v2

# GEO: Generative Engine Optimization

Abstract.
The advent of large language models (LLMs) has ushered in a new paradigm of search engines that use generative models to gather and summarize information to answer user queries. This emerging technology, which we formalize under the unified framework of generative engines (GEs), can generate accurate and personalized responses, rapidly replacing traditional search engines like Google and Bing. Generative Engines typically satisfy queries by synthesizing information from multiple sources and summarizing them using LLMs. While this shift significantly improves user utility and generative search engine traffic, it poses a huge challenge for the third stakeholder – website and content creators. Given the black-box and fast-moving nature of generative engines, content creators have little to no control over when and how their content is displayed. With generative engines here to stay, we must ensure the creator economy is not disadvantaged. To address this, we introduce Generative Engine Optimization (GEO), the first novel paradigm to aid content creators in improving their content visibility in GE responses through a flexible black-box optimization framework for optimizing and defining visibility metrics. We facilitate systematic evaluation by introducing GEO-bench, a large-scale benchmark of diverse user queries across multiple domains, along with relevant web sources to answer these queries. Through rigorous evaluation, we demonstrate that GEO can boost visibility by up to 40% in GE responses. Moreover, we show the efficacy of these strategies varies across domains, underscoring the need for domain-specific optimization methods. Our work opens a new frontier in information discovery systems, with profound implications for both developers of GEs and content creators.111Code and Data available at https://generative-engines.com/GEO/
1. Introduction
The invention of traditional search engines three decades ago revolutionized information access and dissemination globally. While they were powerful and ushered in a host of applications like academic research and e-commerce, they were limited to providing a list of relevant websites for user queries. However, the recent success of large language models (LLMs) has paved the way for better systems like BingChat, Google’s SGE, and perplexity.ai that combine conventional search engines with generative models. We dub these systems generative engines (GE) because they search for information and generate multi-modal responses by using multiple sources. Technically, generative engines (Figure 2) retrieve relevant documents from a database (like the internet) and use large neural models to generate a response grounded on the sources, ensuring attribution and a way for the user to verify the information.
The usefulness of generative engines for developers and users is evident – users access information faster and more accurately, while developers craft precise and personalized responses, improving user satisfaction and revenue. However, generative engines disadvantage the third stakeholder – website and content creators. Generative Engines, in contrast to traditional search engines, remove the need to navigate to websites by directly providing a precise and comprehensive response, potentially reducing organic traffic to websites and impacting their visibility. With millions of small businesses and individuals relying on online traffic and visibility for their livelihood, generative engines will significantly disrupt the creator economy. Further, the black-box and proprietary nature of generative engines makes it difficult for content creators to control and understand how their content is ingested and portrayed.
In this work, we propose the first general creator-centric framework to optimize content for generative engines, which we dub Generative Engine Optimization (GEO), to empower content creators to navigate this new search paradigm. GEO is a flexible black-box optimization framework for optimizing web content visibility for proprietary and closed-source generative engines (Figure 1). GEO ingests a source website and outputs an optimized version by tailoring and calibrating the presentation, text style, and content to increase visibility in generative engines.
Further, GEO introduces a flexible framework for defining visibility metrics tailor-made for generative engines as the notion of visibility in generative engines is more nuanced and multi-faceted than traditional search engines (Figure 3). While average ranking on the response page is a good measure of visibility in traditional search engines, which present a linear list of websites, this does not apply to generative engines. Generative Engines provide rich, structured responses and embed websites as inline citations in the response, often embedding them with different lengths, at varying positions, and with diverse styles. This necessitates the need for visibility metrics tailor-made for generative engines, which measure the visibility of attributed sources over multiple dimensions, such as relevance and influence of citation to query, measured through both an objective and a subjective lens.
To facilitate faithful and extensive evaluation of GEO methods, we propose GEO-bench, a benchmark consisting of 10000 queries from diverse domains and sources, adapted for generative engines. Through systematic evaluation, we demonstrate that our proposed Generative Engine Optimization methods can boost visibility by up to 40% on diverse queries, providing beneficial strategies for content creators. Among other things, we find that including citations, quotations from relevant sources, and statistics can significantly boost source visibility, with an increase of over 40% across various queries. We also demonstrate the efficacy of Generative Engine Optimization on Perplexity.ai, a state-of-the-art real-world generative engine and demonstrate visibility improvements up to 37%.
In summary, our contributions are three-fold:
(1) We propose Generative Engine Optimization, the first general optimization framework for website owners to optimize their websites for generative engines. Generative Engine Optimization can improve the visibility of websites by up to 40% on a wide range of queries, domains, and real-world black-box generative engines.
(2) Our framework proposes a comprehensive set of visibility metrics specifically designed for generative engines and enables content creators to flexibly optimize their content through customized visibility metrics.
(3) To foster faithful evaluation of GEO methods in generative engines, we propose the first large-scale benchmark consisting of diverse search queries from wide-ranging domains and datasets specially tailored for Generative Engines.
2. Formulation & Methodology
2.1. Formulation of Generative Engines
Despite the deployment of numerous generative engines to millions of users, there is currently no standard framework. We provide a formulation that accommodates various modular components in their design. We describe a generative engine, which includes several backend generative models and a search engine for source retrieval. A Generative Engine (GE) takes a user query and returns a natural language response , where represents personalized user information. The GE can be represented as a function:
Generative Engines comprise two crucial components: a.) A set of generative models , each serving a specific purpose like query reformulation or summarization, and b.) A search engine that returns a set of sources given a query . We present a representative workflow in Figure 2, which at the time of writing, closely resembles the design of BingChat. This workflow breaks down the input query into a set of simpler queries that are easier to consume for the search engine. Given a query, a query re-formulating generative model, , generates a set of queries , which are then passed to the search engine to retrieve a set of ranked sources . The sets of sources are passed to a summarizing model , which generates a summary for each source in , resulting in the summary set (). The summary set is passed to a response-generating model , which generates a cumulative response backed by sources . In this work, we focus on single-turn Generative Engines, but the formulation can be extended to multi-turn Conversational Generative Engines (Appendix A).
The response is typically a structured text with embedded citations. Citations are important given the tendency of LLMs to hallucinate information (Ji et al., 2023). Specifically, consider a response composed of sentences . Each sentence may be backed by a set of citations that are part of the retrieved set of documents . An ideal generative engine should ensure all statements in the response are supported by relevant citations (high citation recall), and all citations accurately support the statements they’re associated with (high citation precision) (Liu et al., 2023b). We refer readers to Figure 3 for a representative generative engine response.
2.2. Generative Engine Optimization
The advent of search engines led to search engine optimization (SEO), a process to help website creators optimize their content to improve search engine rankings. Higher rankings correlate with increased visibility and website traffic. However, traditional SEO methods are not directly applicable to Generative Engines. This is because, unlike traditional search engines, the generative model in generative engines is not limited to keyword matching, and the use of language models in ingesting source documents and response generation results in a more nuanced understanding of text documents and user query. With generative engines rapidly emerging as the primary information delivery paradigm and SEO not directly applicable, new techniques are needed. To this end, we propose Generative Engine Optimization, a new paradigm where content creators aim to increase their visibility (or impression) in generative engine responses. We define the visibility of a website (also referred as citation) in a cited response by the function , which the website creator wants to maximize. From the generative engine’s perspective, the goal is to maximize the visibility of citations most relevant to the user query, i.e., maximize ), where measures the relevance of citation to the query in the context of response and is determined by the exact algorithmic design of generative engine and is a black-box function to end-users. Further, both the functions and are subjective and not well-defined yet for generative engines, and we define them next.
2.2.1. Impressions for Generative Engines
In SEO, a website’s impression (or visibility) is determined by its average ranking over a range of queries. However, generative engines’ output nature necessitates different impression metrics. Unlike search engines, Generative Engines combine information from multiple sources in a single response. Factors such as length, uniqueness, and presentation of the cited website determine the true visibility of a citation. Thus, as illustrated in Figure 3, while a simple ranking on the response page serves as an effective metric for impression and visibility in conventional search engines, such metrics are not applicable to generative engine responses.
To address this, we propose several impression metrics. The “Word Count” metric is the normalized word count of sentences related to a citation. Mathematically, this is defined as:
Here is the set of sentences citing , is the set of sentences in the response, and is the number of words in sentence . In cases where a sentence is cited by multiple sources, we share the word count equally with all the citations. Intuitively, a higher word count correlates with the source playing a more important part in the answer, and thus the user gets higher exposure to that source. However, since “Word Count” is not impacted by the ranking of the citations (whether it appears first, for example), we propose a position-adjusted count that reduces the weight by an exponentially decaying function of the citation position:
Intuitively, sentences that appear first in the response are more likely to be read, and the exponent term in definition gives higher weightage to such citations. Thus, a website cited at the top may have a higher impression despite having a lower word count than a website cited in the middle or end of the response. While the above impression metrics are objective and well-grounded, they ignore the subjective aspects of the impact of citations on the user’s attention. To address this, we propose the ”Subjective Impression” metric, which incorporates facets such as the relevance of the cited material to the user query, influence of the citation, uniqueness of the material presented by a citation, subjective position, subjective count, probability of clicking the citation, and diversity in the material presented. We use G-Eval (Liu et al., 2023a), the current state-of-the-art for evaluation with LLMs, to measure each of these sub-metrics.
2.2.2. Generative Engine Optimization methods for website
To improve impression metrics, content creators must make changes to their website content. We present several generative engine-agnostic strategies, referred to as Generative Engine Optimization methods (GEO). Mathematically, every GEO method is a function , where is the initial web content, and is the modified content after applying the GEO method. The modifications can range from simple stylistic alterations to incorporating new content in a structured format. A well-designed GEO is equivalent to a black-box optimization method that, without knowing the exact algorithmic design of generative engines, can increase the website’s visibility and implement textual modifications to independent of the exact queries.
For our experiments, we apply Generative Engine Optimization methods on website content using a large language model, prompted to perform specific stylistic and content changes to the website. In particular, based on the GEO method defining a specific set of desired characteristics, the source content is modified accordingly. We propose and evaluate several such methods:
1: Authoritative: Modifies text style of the source content to be more persuasive and authoritative, 2. Statistics Addition: Modifies content to include quantitative statistics instead of qualitative discussion, wherever possible, 3. Keyword Stuffing: Modifies content to include more keywords from the query, as expected in classical SEO optimization. 4. Cite Sources & 5. Quotation Addition: Adds relevant citations and quotations from credible sources, 6.) 6. Easy-to-Understand: Simplifies the language of website, while 7. Fluency Optimization improves the fluency of website text. 8. Unique Words & 9. Technical Terms: involves adding unique and technical terms respectively wherever possible,
These methods cover diverse general strategies that website owners can implement quickly and use regardless of the website content. Further, except for methods 3, 4, and 5, the remaining methods enhance the presentation of existing content to increase its persuasiveness or appeal to the generative engine, without requiring extra content. On the other hand, methods 3,4, and 5 may require some form of additional content. To analyze the performance gain of our methods, for each input user query, we randomly select one source website to be optimized and apply each of the GEO methods separately on the same source. We refer readers to Appendix B.4 for exact prompts and more details on GEO methods.
3. Experimental Setup
3.1. Evaluated Generative Engine
In accordance with previous works (Liu et al., 2023b), we use a 2-step setup for Generative Engine design. The first step involves fetching relevant sources for input query, followed by a second step where an LLM generates a response based on the fetched sources. Due to context length limitations and quadratic scaling cost based on the context size of transformer models, only the top 5 sources are fetched from the Google search engine for every query. The setup closely mimics the workflow used in previous works and the general design adopted by commercial GEs such as you.com, perplexity.ai. The answer is then generated by gpt3.5-turbo model using the prompt same as prior work (Liu et al., 2023b). We refer readers to Appendix B for more details.
Further in Section C.1, we evaluate the same Generative Engine Optimization methods on Perplexity.ai, which is a commercially deployed generative engine, highlighting the generalizability of our proposed Generative Engine Optimization methods.
3.2. Benchmark : GEO-bench
Since there is currently no publicly available dataset containing Generative Engine related queries, we curate GEO-bench, a benchmark consisting of 10K queries from multiple sources, repurposed for generative engines, along with synthetically generated queries. The benchmark includes queries from nine different sources; each further categorized based on their target domain, difficulty, query intent, and other dimensions.
Datasets:
1. MS Macro, 2. ORCAS-1, and 3. Natural Questions: (Kwiatkowski et al., 2019; Alexander et al., 2022; Craswell et al., 2021) These datasets contain real anonymized user queries from Bing and Google Search Engines. These three collectively represent the common set of datasets that are used in search engine related research. However, Generative Engines will be posed with far more difficult and specific queries with the intent of synthesizing answers from multiple sources instead of searching for them. To this end, we repurpose several other publicly available datasets: 4. AllSouls: This dataset contains essay questions from ”All Souls College, Oxford University”. The queries in this dataset require Generative Engines to perform appropriate reasoning to aggregate information from multiple sources. 5. LIMA: (Zhou et al., 2023) contains challenging questions requiring Generative Engines to not only aggregate information but also perform suitable reasoning to answer the question (eg: writing a short poem, python code.). 6. Davinci-Debtate (Liu et al., 2023b) contains debate questions generated for testing Generative Engines. 7. Perplexity.ai Discover222https://www.perplexity.ai/discover: These queries are sourced from Perplexity.ai’s Discover section, which is an updated list of trending queries on the platform. 8. ELI-5333https://huggingface.co/datasets/eli5_category: This dataset contains questions from the ELI5 subreddit, where users ask complex questions and expect answers in simple, layman’s terms. 9. GPT-4 Generated Queries: To supplement diversity in query distribution, we prompt GPT-4 to generate queries ranging from various domains (eg, science, history) and based on query intent (eg, navigational, transactional) and based on difficulty and scope of generated response (eg, open-ended, fact-based).
Our benchmark comprises 10K queries divided into 8K, 1K, and 1K for train, validation, and test splits, respectively. We preserve the real-world query distribution, with our benchmark containing 80% informational queries and 10% each for transactional and navigational queries. Each query is augmented with the cleaned text content of the top 5 search results from the Google search engine.
Tags
Optimizing website content often requires targeted changes based on the task’s domain. Additionally, a user of Generative Engine Optimization may need to identify an appropriate method for only a subset of queries, considering multiple factors such as domain, user intent, and query nature. To facilitate this, we tag each query with one of seven different categories. For tagging, we employ the GPT-4 model and manually verify high recall and precision on the test split.
Overall, GEO-bench consists of queries from 25 diverse domains such as Arts, Health, and Games; it features a range of query difficulties from simple to multi-faceted; includes 9 different types of queries such as informational and transactional; and encompasses 7 different categorizations. Owing to its specially designed high diversity, the size of the benchmark, and its real-world nature, GEO-bench is a comprehensive benchmark for evaluating Generative Engines and serves as a standard testbed for assessing them for various purposes in this and future works. We provide more details about GEO-bench in Appendix B.2.
3.3. GEO Methods
We evaluate 9 different proposed GEO methods as described in Section 2.2.2. We compare them with a baseline, which measures the impression metric of unmodified website sources. We evaluate methods on the complete GEO-bench test split. Further, to reduce variance in results, we run our experiments on five different random seeds and report the average.
3.4. Evaluation Metrics
We utilize the impression metrics as defined in Section 2.2.1. Specifically, we employ two impression metrics: 1. Position-Adjusted Word Count, which combines word count and position count. To analyze the effect of individual components, we also report scores on the two sub-metrics separately. 2. Subjective Impression, which is a subjective metric encompassing seven different aspects: 1) relevance of the cited sentence to the user query, 2) influence of the citation, assessing the extent to which the generated response relies on the citation, 3) uniqueness of the material presented by a citation, 4) subjective position, gauging the prominence of the positioning of source from the user’s viewpoint, 5) subjective count, measuring the amount of content presented from the citation as perceived by the user, 6) likelihood of the user clicking the citation, and 7) diversity of the material presented. These sub-metrics assess diverse aspects that content creators can target to improve one or more areas effectively. Each sub-metric is evaluated using GPT-3.5, following a methodology akin to that described in G-Eval (Liu et al., 2023a). In G-Eval, a form-based evaluation template is provided to the language model, along with a GE generated response with citations. The model outputs a score (computed by sampling multiple times) for each citation. However, since G-Eval scores are poorly calibrated, we normalize them to have the same mean and variance as Position-Adjusted Word Count to enable a fair and meaningful comparison. We provide the exact templates used in Appendix B.3.
Furthermore, all impression metrics are normalized by multiplying them with a constant factor so that the sum of the impressions of all citations in a response equals 1. In our discussion, we compare methods by calculating the relative improvement in impression. For an initial generated response from sources , and a modified response , the relative improvement in impression for each source is measured as:
The modified response is produced by applying the GEO method being evaluated to one of the sources . The source selected for optimization is chosen randomly but remains constant for a particular query across all GEO methods.
| Method | Position-Adjusted Word Count | Subjective Impression | |||||||||
| Word | Position | Overall | Rel. | Infl. | Unique | Div. | FollowUp | Pos. | Count | Average | |
| Performance without Generative Engine Optimization | |||||||||||
| No Optimization | |||||||||||
| Non-Performing Generative Engine Optimization methods | |||||||||||
| Keyword Stuffing | |||||||||||
| Unique Words | |||||||||||
| High-Performing Generative Engine Optimization methods | |||||||||||
| Easy-to-Understand | |||||||||||
| Authoritative | |||||||||||
| Technical Terms | |||||||||||
| Fluency Optimization | |||||||||||
| Cite Sources | |||||||||||
| Quotation Addition | 27.2 | 24.7 | |||||||||
| Statistics Addition | |||||||||||
4. Results
We evaluate various Generative Engine Optimization methods designed to optimize website content for better visibility in Generative Engine responses, compared against a baseline with no optimization. Our evaluation used GEO-bench, a diverse benchmark of user queries from multiple domains and settings. Performance was measured using two metrics: Position-Adjusted Word Count and Subjective Impression. The former consider word count and citation position in the GE’s response, while the latter computes multiple subjective factors, giving an overall impression score.
Table 1 details the absolute impression metrics of different methods on multiple metrics. The results reveal that our GEO methods consistently outperform the baseline across all metrics on GEO-bench. This shows the robustness of these methods to varying queries, yielding significant improvements despite query diversity. Specifically, our top-performing methods, Cite Sources, Quotation Addition, and Statistics Addition, achieved a relative improvement of 30-40% on the Position-Adjusted Word Count metric and 15-30% on the Subjective Impression metric. These methods, involving adding relevant statistics (Statistics Addition), incorporating credible quotes (Quotation Addition), and including citations from reliable sources (Cite Sources) in the website content, require minimal changes but significantly improve visibility in GE responses, enhancing both the credibility and richness of the content.
Interestingly, stylistic changes such as improving fluency and readability of the source text (Fluency Optimization and Easy-to-Understand) also resulted in a significant visibility boost of 15-30%. This suggests that Generative Engines value not only content but also information presentation.
Further, given generative models are often designed to follow instructions, one would expect a more persuasive and authoritative tone in website content to boost visibility. However, we find no significant improvement, demonstrating that Generative Engines are already somewhat robust to such changes. This highlights the need for website owners to focus on improving content presentation and credibility.
Finally, we evaluate keyword stuffing, i.e., adding more relevant keywords to website content. While widely used for Search Engine Optimization, we find such methods offer little to no improvement on generative engine’s responses. This underscores the need for website owners to rethink optimization strategies for generative engines, as techniques effective in search engines may not translate to success in this new paradigm.
5. Analysis
| Method | Relative Improvement (%) in Visibility | ||||
| Rank-1 | Rank-2 | Rank-3 | Rank-4 | Rank-5 | |
| Authoritative | -6.0 | 4.1 | -0.6 | 12.6 | 6.1 |
| Fluency Opt. | -2.0 | 5.2 | 3.6 | -4.4 | 2.2 |
| Cite Sources | -30.3 | 2.5 | 20.4 | 15.5 | 115.1 |
| Quotation Addition | -22.9 | -7.0 | 3.5 | 25.1 | 99.7 |
| Statistics Addition | -20.6 | -3.9 | 8.1 | 10.0 | 97.9 |
| Method | Top Performing Tags | ||
| Rank-1 | Rank-2 | Rank-3 | |
| Authoritative | Debate | History | Science |
| Fluency Opt. | Business | Science | Health |
| Cite Sources | Statement | Facts | Law & Gov. |
| Quotation Addition | People & Society | Explanation | History |
| Statistics Addition | Law & Gov. | Debate | Opinion |
5.1. Domain-Specific Generative Engine Optimizations
In Section 4, we presented the improvements achieved by GEO across the entirety of the GEO-bench benchmark. However, in real-world SEO scenarios, domain-specific optimizations are often applied. With this in mind, and considering that we provide categories for every query in GEO-bench, we delve deeper into the performance of various GEO methods across these categories.
Table 3 provides a detailed breakdown of the categories where our GEO methods have proven to be most effective. A careful analysis of these results reveals several intriguing observations. For instance, Authoritative significantly improves performance in debate-style questions and queries related to the “historical” domain. This aligns with our intuition, as a more persuasive form of writing is likely to hold more value in debates.
Similarly, the addition of citations through Cite Sources is particularly beneficial for factual questions, likely because citations provide a source of verification for the facts presented, thereby enhancing the credibility of the response. The effectiveness of different GEO methods varies across domains. For example, as shown in row 5 of Table 3, domains such as ‘Law & Government’ and question types like ‘Opinion’ benefit significantly from the addition of relevant statistics in the website content, as implemented by Statistics Addition. This suggests that data-driven evidence can enhance the visibility of a website in particular contexts. The method Quotation Addition is most effective in the ‘People & Society’, ‘Explanation’, and ‘History’ domains. This could be because these domains often involve personal narratives or historical events, where direct quotes can add authenticity and depth to the content. Overall, our analysis suggests that website owners should strive towards making domain-specific targeted adjustments to their websites for higher visibility.
5.2. Optimization of Multiple Websites
In the evolving landscape of Generative Engines, GEO methods are expected to become widely adopted, leading to a scenario where all source contents are optimized using GEO. To understand the implications, we conducted an evaluation of GEO methods by optimizing all source contents simultaneously, with results presented in Table 3. A key observation is the differential impact of GEO on websites based on their Search Engine Results Pages (SERP) ranking. Notably, lower-ranked websites, which typically struggle for visibility, benefit significantly more from GEO. This is because traditional search engines rely on multiple factors, such as the number of backlinks and domain presence, which are challenging for small creators to achieve. However, since Generative Engines utilize generative models conditioned on website content, factors such as backlink building should not disadvantage small creators. This is evident from the relative improvements in visibility shown in Table 3. For example, the Cite Sources method led to a substantial 115.1% increase in visibility for websites ranked fifth in SERP, while on average, the visibility of the top-ranked website decreased by 30.3%.
This finding highlights GEO’s potential as a tool to democratize the digital space. Many lower-ranked websites are created by small content creators or independent businesses, who traditionally struggle to compete with larger corporations in top search engine results. The advent of Generative Engines might initially seem disadvantageous to these smaller entities. However, the application of GEO methods presents an opportunity for these content creators to significantly improve their visibility in Generative Engine responses. By enhancing their content with GEO, they can reach a wider audience, leveling the playing field and allowing them to compete more effectively with larger corporations.
5.3. Combination of GEO Strategies
While individual GEO strategies show significant improvements across various domains, in practice, website owners are expected to employ multiple strategies in conjunction. To study the performance improvements achieved by combining GEO strategies, we consider all pairs of combinations of the top 4 performing GEO methods, namely Cite Sources, Fluency Optimization, Statistics Addition, and Quotation Addition. Figure 4 displays the heatmap of relative improvement in the Position-Adjusted Word Count visibility metric achieved by combining different GEO strategies. The analysis demonstrates that the combination of Generative Engine Optimization methods can enhance performance, with the best combination (Fluency Optimization and Statistics Addition) outperforming any single GEO strategy by more than 5.5%444Due to cost constraints, the analysis was conducted on a subset of 200 examples from the test split, and therefore the numbers presented here differ from those in Table 1. Furthermore, Cite Sources significantly boosts performance when used in conjunction with other methods (Average: 31.4%), despite it being relatively less effective when used alone (8% lower than Quotation Addition). The findings underscore the importance of studying GEO methods in combination, as they are likely to be used by content creators in the real world.
5.4. Qualitative Analysis
We present a qualitative analysis of GEO methods in Table 4, containing representative examples where GEO methods boost source visibility with minimal changes. Each method optimizes a source through suitable text additions and deletions. In the first example, we see that simply adding the source of a statement can significantly boost visibility in the final answer, requiring minimal effort from the content creator. The second example demonstrates that adding relevant statistics wherever possible ensures increased source visibility in the final Generative Engine response. Finally, the third row suggests that merely emphasizing parts of the text and using a persuasive text style can also lead to improvements in visibility.
| Method | GEO Optimization | Relative Improvement |
| Query: What is the secret of Swiss chocolate | ||
| With per capita annual consumption averaging between 11 and 12 kilos, Swiss people rank among the top chocolate lovers in the world (According to a survey conducted by The International Chocolate Consumption Research Group [1]) | ||
| Cite Sources | 132.4% | |
| Query: Should robots replace humans in the workforce? | ||
| Source: Not here, and not now — until recently. The big difference is that robots have come not to destroy our lives, but to disrupt our work, with a staggering 70% increase in robotic involvement in last decade. | ||
| Statistics Addition | 65.5% | |
| Query: Did the jacksonville jaguars ever make it to the superbowl? | ||
| Source: It is important to note that The Jaguars have never appeared made an appearance in the Super Bowl. However, They have achieved an impressive feat by securing 4 divisional titles to their name. , a testament to their prowess and determination. | ||
| Authoritative | 89.1% |
6. GEO in the Wild : Experiments with Deployed Generative Engine
| Method | Position-Adjusted Word Count | Subjective Impression |
| No Optimization | ||
| Keyword Stuffing | ||
| Quotation Addition | 29.1 | |
| Statistics Addition | 33.9 |
To reinforce the efficacy of our proposed Generative Engine Optimization methods, we evaluate them on Perplexity.ai, a real deployed Generative Engine with a large user base. Results are in Table 5. Similar to our generative engine, Quotation Addition performs best in Position-Adjusted Word Count with a 22% improvement over the baseline. Methods that performed well in our generative engine such as Cite Sources, Statistics Addition show improvements of up to 9% and 37% on the two metrics. Our observations, such as the ineffectiveness of traditional SEO methods like Keyword Stuffing, are further highlighted, as it performs 10% worse than the baseline. The results are significant for three reasons: 1) they underscore the importance of developing different Generative Engine Optimization methods to benefit content creators, 2) they highlight the generalizability of our proposed GEO methods on different generative engines, 3) they demonstrate that content creators can use our easy-to-implement proposed GEO methods directly, thus having a high real-world impact. We refer readers to Appendix C.1 for more details.
7. Related Work
Evidence-based Answer Generation:
Previous works have used several techniques for answer generation backed by sources. Nakano et al. (2021) trained GPT-3 to navigate web environments to generate source-backed answers. Similarly, other methods (Shuster et al., 2022; Thoppilan et al., 2022; Menick et al., 2022) fetch sources via search engines for answer generation. Our work unifies these approaches and provide a common benchmark for improving these systems in the future. In a recent working draft, Kumar and Lakkaraju (Kumar and Lakkaraju, 2024) showed that strategic text sequences can manipulate LLM recommendations to enhance product visibility in generative engines. While their approach focuses on increasing product visibility through adversarial text, our method introduces non-adversarial strategies to optimize any website content for improved visibility in generative engine search results.
Retrieval-Augmented Language Models:
Several recent works have tackled the issues of limited memory of language models by fetching relevant sources from a knowledge base to complete a task (Asai et al., 2021; Mialon et al., 2023; Guu et al., 2020). However, Generative Engine needs to generate an answer and provide attributions throughout the answer. Further, Generative Engine is not limited to a single text modality regarding both input and output. Additionally, the framework of Generative Engine is not limited to fetching relevant sources but instead comprises multiple tasks such as query reformulation, source selection, and making decisions on how and when to perform them.
Search Engine Optimization:
In nearly the past 25 years, extensive research has optimized web content for search engines (Ankalkoti, 2017; Shahzad et al., 2020; Kumar et al., 2019). These methods fall into On-Page SEO, improving content and user experience, and Off-Page SEO, boosting website authority through link building. In contrast, GEO deals with a more complex environment involving multi-modality, conversational settings. Since GEO is optimized against a generative model not limited to simple keyword matching, traditional SEO strategies will not apply to Generative Engine settings, highlighting the need for GEO.
8. Conclusion
In this work, we formulate search engines augmented with generative models that we dub generative engines. We propose Generative Engine Optimization (GEO) to empower content creators to optimize their content under generative engines. We define impression metrics for generative engines and propose and release GEO-bench: a benchmark encompassing diverse user queries from multiple domains and settings, along with relevant sources needed to answer those queries. We propose several ways to optimize content for generative engines and demonstrate that these methods are capable of boosting source visibility by up to 40% in generative engine responses. Among other findings, we show that including citations, quotations from relevant sources, and statistics can significantly boost source visibility. Further, we discover a dependence of the effectiveness of GEO methods on the domain of the query, and potential of combining multiple GEO strategies in conjunction. We show promising results on a commercially deployed generative engine with millions of active users showcasing the real-world impact of our work. Our work serves as a first step towards understanding the impact of generative engines on the digital space and the role of GEO in this new paradigm of search engines.
9. Limitations
While we rigorously test our proposed methods on two generative engines, including a publicly available one, methods may need to adapt over time as GEs evolve, mirroring the evolution of SEO. Additionally, despite our efforts to ensure the queries in our GEO-bench closely resemble real-world queries, the nature of queries can change over time, necessitating continuous updates. Lastly, while every query in our proposed GEO-bench is tagged and manually inspected, there may be discrepancies due to subjective interpretations or errors in labeling.
Appendix A Conversational Generative Engine
In Section 2.1, we discussed a single-turn Generative Enginethat outputs a single response given the user query. However, one of the strengths of upcoming Generative Engines will be their ability to engage in an active back-and-forth conversation with the user. The conversation allows users to provide clarifications to their queries or Generative Engine response and ask follow-ups. Specifically, in equation 1, instead of the input being a single query , it is modeled as a conversation history pairs. The response is then defined as:
where is the turn number.
Further, to engage the user in a conversation, a separate LLM, or , may generate suggested follow-up queries based on , , and . The suggested follow-up queries are typically designed to maximize the likelihood of user engagement. This not only benefits Generative Engine providers by increasing user interaction, but also benefits website owners by enhancing their visibility. Furthermore, these follow-up queries can help users by getting more detailed information.
Appendix B Experimental Setup
B.1. Evaluated Generative Engine
While the exact design specifications of popular Generative Engines are not public, based on previous works (Liu et al., 2023b), and our experiments, most of them follow 2-step procedure for generating responses. The first step involves fetching relevant sources for input query, followed by a LLM generating response based on the fetched sources. In our setup, we fetch top 5 sources from Google search engine for every query. The answer is generated by gpt3.5-turbo model using a prompt same as prior work (Liu et al., 2023b). The model is prompted to output an appropriate response for the query, with each sentence cited by one of the 5 sources provided. We sample 5 different answers at temperature=0.7 and top_p=1, to reduce statistical deviations. The exact prompt used is shown in Listing 1.
B.2. Benchmark
Since, currently there is no publicly available dataset containing Generative Engine related queries we curate GEO-bench, a benchmark containing 10K queries from multiple sources repurposed for generative engines along with synthetically generated queries. The benchmark contains queries from nine different sources, each further categorized based on their target domain, difficulty, query intent and other dimensions. The datasets used in constructing the benchmark are: 1. MS Macro & 2. ORCAS-1: contains real anonymized user queries from Bing Search Engine. and 3. Natural Questions: containing queries from Google Search Engine. These three collectively represent the common set of datasets that are used in search engine related research. However, Generative Engines will be posed with far more dofficult and specific queries with intent of synthesizing answer from multiple sources instead of search for them. To this end, we re-purpose several other publicly available datasets: 4. AllSouls: A dataset containing essay questions from “All Souls College, Oxford University”. The queries in the dataset cannot be usually answered from a single source, and requires Generative Engines to aggregate information multiple sources and perform reasonable reasoning on them. 5. LIMA (Zhou et al., 2023) contains carefully crafted queries and responses for training pretrained language models for instruction following. The queries in the dataset represent a more challenging distribution of queries asked in Generative Engine, and often requires LLM’s creative and technical powress to generate answers (eg: writing short poem, python code.) 6. Perplexity.ai Discover: These queries are sourced from Perplexity.ai’s, a public Generative Engine, Discover section which is a updated list of trending queries on the platform. These queries represent a real distribution of queries made on Generative Engines. 7. Davinci-Debtate (Liu et al., 2023b) contains debate questions generated using text-davinci-003 and sourced from Perspectrum dataset (Chen et al., 2019). This dataset were specifically designed for Generative Engines. 8. ELI-5: contains questions from the ELI5 subreddit, where users ask complex questions and expect answers in simple, layman terms. 9. GPT-4 Generated Queries: To further supplement diversity in query distribution and increase Generative Engine specific queries, we prompt GPT-4 to generate queries ranging from various domains (eg: science, history), based on query intent (eg: navigational, transactional), based on difficulty and scope of generated response (eg: open-ended, fact-based). Representative queries from each of the datasets are shown in Figure 2.
In total our benchmark contains 10K queries split into 8K,1K,1K train/val/test splits. Every query is tagged into multiple categories gauging various dimensions such as intent, difficulty, domain of query and format of answer type using GPT-4. In terms of query intent, we maintain the real-world query distribution, with our benchmark containing 80% informational queries, and 10% transactional queries and 10% navigational queries (Jansen et al., 2008). Further, we augment every query with cleaned text content of top 5 search results from Google search engine. Owing to specially designed high benchmark diversity, size, complexity and real-world nature, we believe GEO-bench is a comprehensive benchmark for evaluating Generative Engines and serves as a standard testbed for evaluating Generative Engines for multiple purposes in this and future works.
Tags:
Optimizing website content often requires making targeted changes based on the domain of the task. Further, a user of Generative Engine Optimization may need to find an appropriate method for only a subset of queries based on multiple factors, such as domain, user intent, query nature. To this end, we tag each of the queries based from a pool of 7 different categories. For tagging we use GPT-4 model, and manually confirm high recall and precision in tagging. However, owing to such automated system, the tags can be noisy and should not be considered. Detail about each of these queries are presented here and Listing 5
Difficulty Level: The complexity of the query, ranging from simple to complex.
1# Example of a simple query2query = "What is the capital of France?"3# Example of a complex query4query = "What are the implications of the Schrodinger equation in quantum mechanics?" -
Nature of Query: The type of information sought by the query, such as factual, opinion, or comparison.
1# Example of a factual query2query = "How does a car engine work?"3# Example of an opinion query4query = "What is your opinion on the Harry Potter series?" -
Genre: The category or domain of the query, such as arts and entertainment, finance, or science.
1# Example of a query in the arts and entertainment genre2query = "Who won the Oscar for Best Picture in 2020?"3# Example of a query in the finance genre4query = "What is the current exchange rate between the Euro and the US Dollar?" -
Specific Topics: The specific subject matter of the query, such as physics, economics, or computer science.
1# Example of a query on a specific topic in physics2query = "What is the theory of relativity?"3# Example of a query on a specific topic in economics4query = "What is the law of supply and demand?" -
Sensitivity: Whether the query involves sensitive topics or not.
1# Example of a non-sensitive query2query = "What is the tallest mountain in the world?"3# Example of a sensitive query4query = "What is the current political situation in North Korea?" -
User Intent: The purpose behind the user’s query, such as research, purchase, or entertainment.
1# Example of a research intent query2query = "What are the health benefits of a vegetarian diet?"3# Example of a purchase intent query4query = "Where can I buy the latest iPhone?" -
Answer Type: The format of the answer that the query is seeking, such as fact, opinion, or list.
1# Example of a fact answer type query2query = "What is the population of New York City?"3# Example of an opinion answer type query4query = "Is it better to buy or rent a house?"
B.3. Evaluation Metrics
We evaluate all methods by measuring the Relative Improvement in Impression. Specifically, given an initial generated Response from sources s, and a modified response from sources s, we measure the relative improvement in impression of each of the source as:
We use the impression metrics as defined in Section 2.2.1. Specifically, we use two impression metrics: 1. Position-Adjusted Word Count which is a combination of word count and position count. To dissect effect of individual components, we also report individual scores on the 2 sub-metrics. 2. Subjective Impression which is a subjective impression metric which is a combination of seven different aspects: Relevance of citation to query, influence of citaiton on response, diversity and uniqueness of information presented, likelihood of followup by the user, perceived rank and amount of information presented in the answer. All these sub-metrics are evaluated using GPT-3.5, using methodology similar to described as in G-Eval (Liu et al., 2023a). However, since G-Eval scores are ill-calibrated, we need to suitably normalize them for fair and appropriate comparison. We normalize Subjective Impression scores with respect to baseline scores of Position-Adjusted Word Count to ensure same mean and standard deviation. We list prompts for each of the sub-metrics in Listing 6 7 8 9.
B.4. GEO Methods
To improve the impression metrics, content creators need to make changes to their websites. To this end, we present several generative engine-agnostic strategies, refered to as Generative Engine Optimization methods (GEO). Mathematically, every GEO method is a function , where is the initial web content, and is the modified website content after applying LEO method. We propose and evaluate a series of methods.
1: Authoritative: Modifies text style of the source content to be more persuasive while making authoritative claims, 2. Statistics Addition: Modifies content to include quantitative statistics instead of qualitative discussion, wherever possible, 3. Keyword Stuffing: Modifies content to include more keywords from the query, as would be expected in classical SEO optimization. 4. Cite Sources & 5. Quotation Addition: Adds relevant citations and quotations from credible sources, 6.) 6. Easy-to-Understand: Simplifies the language of website, while 7. Fluency Optimization improves the fluency of website text. 8. Unique Words & 9. Technical Terms: involves adding unique and technical terms respectively wherever posssible,
In order to analyse the performance gain of our methods, for each input query, we randomly select one source to be optimized using all GEO separately. Further, for every method, 5 answers are generated per query to reduce statistical noise in the results.
B.5. Prompts for GEO methods
Here we present prompts used for different Generative Engine Optimization methods. GPT-3.5 turbo was used for all experiments.
Appendix C Results
We perform experiments on 5 seeds, and present results in Table 6
C.1. GEO in the Wild : Experiments with Deployed Generative Engine
To further reinforce the efficacy of our proposed Generative Engine Optimization methods, we evaluate them on Perplexity.ai a deployed Generative Engine with a large user base. Since perplexity.ai does not allow the user to specify source URLs, we instead provide source text as file uploads to perlexity.ai. We ensure all answers are generated only using the file sources provided. We evaluate all our methods on a subset of 200 samples of our test set. Other experimental procedures are the same as our main results. Results using Perplexity.ai are shown in Table 7. We find, similar to our generative engine Quotation Addition performs the best in Position-Adjusted Word Count with a relative improvement of 22% over the baseline. Further, methods that performed well in our generative engine such as Cite Sources, Statistics Addition show high improvements of up to 9% and 37% on the two metrics. Further, our observations such as the ineffectiveness of traditional methods used in SEO such as Keyword Stuffing are further highlighted, as it performs 10% worse than the baseline. The result underscores the importance of developing different Generative Engine Optimization methods to benefit the content-creators and further highlights that our simple-to-implement proposed methods can be used directly by content-creators, thus having a high real-world impact.
Appendix D Ethical Considerations and Reproducibility Statement
In our study, we focus on enhancing the visibility of websites in generative engines. We do not directly interact with sensitive data or individuals. While the sources we retrieve from search engines may contain biased or inappropriate content, these are already publicly accessible, and our study neither amplifies nor endorses such content. We believe that our work is ethically sound as it primarily deals with publicly available information and aims to improve the user experience in generative engines.
Regarding reproducibility, we have made our code available to allow others to replicate our results. Our main experiments have been conducted with five different seeds to minimize potential statistical deviations.
| Method | Position-Adjusted Word Count | Subjective Impression | |||||||||
| Word | Position | Overall | Rel. | Infl. | Unique | Div. | FollowUp | Pos. | Count | Average | |
| Performance without Generative Engine Optimization | |||||||||||
| No Optimization | |||||||||||
| Non-Performing Generative Engine Optimization methods | |||||||||||
| Keyword Stuffing | |||||||||||
| Unique Words | |||||||||||
| High-Performing Generative Engine Optimization methods | |||||||||||
| Easy-to-Understand | |||||||||||
| Authoritative | |||||||||||
| Technical Terms | |||||||||||
| Fluency Optimization | |||||||||||
| Cite Sources | |||||||||||
| Quotation Addition | |||||||||||
| Statistics Addition | |||||||||||
| Method | Position-Adjusted Word Count | Subjective Impression | |||||||||
| Word | Position | Overall | Rel. | Infl. | Unique | Div. | FollowUp | Pos. | Count | Average | |
| Performance without Generative Engine Optimization | |||||||||||
| No Optimization | |||||||||||
| Non-Performing Generative Engine Optimization methods | |||||||||||
| Keyword Stuffing | |||||||||||
| Unique Words | |||||||||||
| High-Performing Generative Engine Optimization methods | |||||||||||
| Authoritative | |||||||||||
| Fluency Optimization | |||||||||||
| Cite Sources | |||||||||||
| Quotation Addition | 29.1 | ||||||||||
| Statistics Addition | 33.9 | ||||||||||